# Kullback–Leibler Potential for Non-Ergodic Replication Dynamics: An Information-Theoretic Second Law

### v1

@@ Abstract
This study aims to quantify and visualize the degradation of fidelity (information degradation) that inevitably accompanies the replication of information within the framework of information thermodynamics and to propose an information-theoretic formulation of the second law based on this phenomenon. While previous research in information thermodynamics has focused on the thermodynamic costs associated with information “erasure” or “measurement” through concepts such as Landauer’s principle and mutual information, little systematic discussion has addressed the inherently irreversible nature of “replication” itself and the accompanying degradation of information structure. In this study, we construct a mathematical model of information replication using a discrete Markov model and Gaussian convolution, and quantify changes in information at each replication step: Shannon entropy, cross-entropy, and the Kullback–Leibler divergence (KLD). The monotonic decrease of KLD exhibits a Lyapunov-like property, which can be interpreted as a potential analogous to the free energy in the process by which a nonequilibrium system converges to a particular steady state. Furthermore, we extend this framework to the potential applicability to biological information processes such as DNA replication, showing that the free energy required for degradation and repair can be expressed in terms of KLD. This contributes to building a unified information-thermodynamic framework for operations such as replication, transmission, and repair of information.
I Introduction
The second law of thermodynamics links irreversibility to entropy production. In information thermodynamics, the physical cost of manipulating information has been clarified for erasure and measurement through Landauer’s principle and its extensions: erasing one bit dissipates at least $k_{\mathrm{B}}T\ln 2$ of heat [1], and feedback/measurement operations admit refined second-law-like inequalities expressed through mutual information [2, 3, 4]. In contrast, the inherently irreversible nature of replication - biological or digital - has not been formulated in a unified information–thermodynamic framework. From a biochemical point of view, kinetic proofreading already indicates that high-fidelity copying requires free energy expenditure [5], yet a coarse-grained information-theoretic law for replication itself has remained elusive.
This work. We model replication as a discrete-time Markov map on probability distributions under block invariance (non-ergodic constraints). Given an initial distribution $p_{0}$, let $\Pi(p_{0})$ denote the reachable steady set composed of blockwise invariant distributions that preserve the initial block masses (formal definitions in Sect. II). Our central object is Kullback–Leibler (KL) potential.
$V(p)\ :=\ \inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}\!\bigl{(}p\,\|\,\pi\bigr{)},$ (1)
which measures the excess information retained beyond the conserved coarse variables. We prove that $V$ is a Lyapunov function for block-invariant replication dynamics $p_{n+1}=Tp_{n}$:
$V(p_{n+1})\ \leq\ V(p_{n})\qquad(n=0,1,2,\dots),$ (2)
with equality only in blockwise steady states (Theorem 1). Moreover, when each block is ergodic with a unique invariant $\pi_{j}^{\ast}$, the potential vanishes asymptotically:
$V(p_{n})\ \xrightarrow[n\to\infty]{}\ 0.$ (3)
Unlike the classical ergodic case—where the KL divergence to a single invariant measure is Lyapunov—our $V$ in (1) is a minimum over steady mixtures constrained by the initial block masses, capturing the essential non-ergodicity of replication.
Instantiations and readouts. We instantiate the framework in two settings: (i) image copying via Gaussian convolution (an information-losing map), and (ii) a block diagonal substitution model for DNA replication with proofreading. In both cases, we report the stepwise evolution of Shannon entropy $H$, cross-entropy $H_{\times}$, KL divergence $D_{\mathrm{KL}}$, and the potential $V$ using the single set of definitions and Theorem 1 from Sect. II. All entropies/divergences are in nats unless noted. Connections to prior information-thermodynamic perspectives on replication and non-equilibrium free energy analogies (e.g., interpreting cross-entropy differences as generated heat) are discussed in the context of our results [19, 20].
II General framework and main theorem
II.1 Block invariance and reachable steady set
The information replication operations we study (image blurring, base substitution with proofreading, etc.) are inherently local: one step only rearranges probability mass within a constrained neighborhood (spatial or compositional), but does not freely mix all microscopic states. As a result, certain coarse observables, e.g., the total mass inside a spatial region of an image, or the adenine(A) -thymine (T) rich block and guanine(G) -cytosine (C) rich block in DNA to be replicated as a template, are effectively conserved by every replication step. When such conserved coarse variables exist, the Markov dynamics does not converge to a single global invariant distribution; instead, the state space decomposes into components that evolve independently. We call this situation non-ergodic in the sense that long-term behavior retains memory of the initial coarse structure through the conserved weights.
@@ II.1 Block invariance and reachable steady set
(i) In the image model, a blockwise Gaussian convolution forbids smoothing across block boundaries; the total intensity in each block is preserved at every step. (ii) In the DNA model, a block diagonal substitution kernel reflects biochemical or sequence constraints (e.g., AT- vs. GC-rich segments); block masses (AT vs. GC) remain fixed while within-block compositions relax. (iii) More generally, locality, topological separation, or routing restrictions in copying channels induce invariant ’blocks’ that prevent global mixing. Because replication preserves these coarse variables, the appropriate equilibrium notion is not a single point but a set of steady states obtained by fully relaxing within each invariant component while keeping the initial coarse weights fixed. This motivates modeling the state space as a disjoint union of blocks and defining a reachable steady set against which we will measure the distance using a KL-based potential. We consider a finite state space partitioned into disjoint blocks.
@@
and a Markov kernel $T$ acting on $\mathcal{P}(\mathcal{X})$. Here $\mathcal{P}(\mathcal{X})$ denotes the probability simplex over $\mathcal{X}$, i.e. the set of all probability distributions $p:\mathcal{X}\to[0,1]$ with $\sum_{x\in\mathcal{X}}p(x)=1$; we identify $p$ with a row vector and let $T$ act by $(pT)(y)=\sum_{x\in\mathcal{X}}p(x)\,T(x,y)$, where for each $x$ the map $A\mapsto T(x,A)$ is a probability measure on subsets of $\mathcal{X}$ (so, in the finite case, $T$ is a row stochastic matrix). The dynamics is block-invariant if the transitions never cross the block boundaries, i.e.,
$x\in\mathcal{X}_{j}\ \Longrightarrow\ T(x,\mathcal{X}_{j})=1\qquad(j=1,\dots,m).$ (5)
In above, for a Markov kernel $T$, $T(x,A)$ is the one-step probability to land in $A$ given the current state $x$. Thus, for $x\in\mathcal{X}_{j}$, the condition $T(x,\mathcal{X}_{j})=1$ means the next state lies in the same block with probability 1; equivalently $\sum_{y\in\mathcal{X}_{j}}T(x,y)=1$ and $T(x,y)=0$ for all $y\notin\mathcal{X}_{j}$. In matrix form, each row indexed by $x\in\mathcal{X}_{j}$ has the support contained in $\mathcal{X}_{j}$, so $T$ is block-diagonal and block masses are preserved.Block invariance guarantees conservation of block masses $w_{j}$ but not convergence within blocks. Our potential $V(p)=\inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p\|\pi)$ over the reachable steady set $\Pi(p_{0})$ is designed to handle precisely these non-ergodic situations: by block decomposition and the data-processing inequality, $V(p_{n+1})\leq V(p_{n})$ always holds; and $V(p_{n})\to 0$ requires, in addition, that each block be primitive (irreducible and aperiodic), yielding a unique invariant $\pi_{j}^{\ast}$. For $p\in\mathcal{P}(\mathcal{X})$, define the block masses
$w_{j}(p)\ :=\ p(\mathcal{X}_{j}),\qquad j=1,\dots,m,$ (6)
and, whenever $w_{j}(p)>0$, the conditional distributions within each block by
$p^{(j)}(x)\ :=\ \begin{cases}\displaystyle\frac{p(x)}{w_{j}(p)},&x\in\mathcal{X}_{j},\\[4.30554pt] 0,&x\notin\mathcal{X}_{j}.\end{cases}$ (7)
Let $T_{j}$ denote the restriction of $T$ to block $\mathcal{X}_{j}$. The set of $T_{j}$-invariant measures is
@@
Given an initial distribution $p_{0}$, reachable steady set is defined as
$\Pi(p_{0})\ :=\ \Bigl{\{}\ \pi=\sum_{j=1}^{m}w_{j}(p_{0})\,\pi_{j}\ :\ \pi_{j}\in\mathcal{I}_{j}\ \Bigr{\}}.$ (9)
Intuitively, $\Pi(p_{0})$ collects all steady states that (i) fully relax within each block and (ii) preserve the initial coarse variables $\{w_{j}(p_{0})\}_{j=1}^{m}$ fixed by non-ergodic constraints.
II.2 KLD potential and two basic lemmas
We adopt the Kullback–Leibler divergence (KLD) between $p,q\in\mathcal{P}(\mathcal{X})$,
$D_{\mathrm{KL}}(p\|q)\ :=\ \sum_{x\in\mathcal{X}}p(x)\,\log\!\frac{p(x)}{q(x)},$ (10)
with the convention $0\log(0/q)=0$ and $p\log(p/0)=+\infty$. Given the reachable steady set $\Pi(p_{0})$ in (9), we define KLD potential
$V(p)\ :=\ \inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p\|\pi),$ (11)
(cf. the introductory statement (1)). The next two lemmas formalize (i) conservation of block masses and conditional evolution, and (ii) blockwise decomposition of KLD.
@@ II.2 KLD potential and two basic lemmas
$(Tp)^{(j)}(A)=\frac{(Tp)(A)}{w_{j}(p_{+})}=\frac{\sum_{x\in\mathcal{X}}p(x)T(x,A)}{w_{j}(p)}\\ =\sum_{x\in\mathcal{X}_{j}}\tfrac{p(x)}{w_{j}(p)}\,T_{j}(x,A)=(T_{j}p^{(j)})(A).$ (14)
@@
Let $\pi=\sum_{j=1}^{m}w_{j}\,\pi_{j}$ with $w_{j}>0$ and $\pi_{j}\in\mathcal{P}(\mathcal{X}_{j})$ supported on disjoint blocks $\mathcal{X}_{j}$. Then, for any $p\in\mathcal{P}(\mathcal{X})$,
@@
Equations (12)–(13) show that the dynamics preserves coarse variables $\{w_{j}\}$ and evolves conditionals within each block, while (15) cleanly separates within-block divergences from a coarse-mass mismatch term. These identities will be the backbone for the Lyapunov property established in Sec. II.3.
II.3 Lyapunov property of the KLD potential
Motivation and novelty. In many replication scenarios the dynamics is inherently local and respects coarse constraints that are fixed by the substrate: for DNA, heterogeneous base composition (e.g., AT/GC content and sequence-context effects) induces blocks within which substitutions and proofreading occur, while inter-block flow is strongly suppressed. Our novelty is to adopt this block structure explicitly and measure nonequilibrium distance not to a single invariant law, but to the reachable steady set that preserves the initial block masses. Concretely, with the block partition and the reachable set $\Pi(p_{0})$ defined in Sec. II.1, we use the KL potential
$V(p)\ :=\ \inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p\|\pi),$
which collapses to the classical KL-to-steady Lyapunov function in the ergodic/single-invariant limit, but remains meaningful under non-ergodicity caused by block invariance.
Statement. We show that $V$ decreases monotonically along any block-invariant Markov evolution $p_{n+1}=Tp_{n}$.
@@ II.3 Lyapunov property of the KLD potential
Let $T$ satisfy the block-invariance condition (5) and define $V$ as above. Then, for all $n\geq 0$,
@@
Fix any $\pi\in\Pi(p_{0})$ and write $\pi=\sum_{j=1}^{m}w_{j}(p_{0})\,\pi_{j}$ with $\pi_{j}\in\mathcal{I}_{j}$. By Lemma 1, $w_{j}(Tp)=w_{j}(p_{0})$ and, by Lemma 2,
@@
For KLD, the data-processing inequality (DPI) for $f$-divergences gives
@@
so inserting (20) into (19) yields
@@
Taking the infimum over $\pi\in\Pi(p_{0})$ on both sides proves (17). If each $T_{j}$ is primitive with invariant $\pi_{j}^{\ast}$, then $p_{n}^{(j)}\to\pi_{j}^{\ast}$, and hence (18). ∎
The DPI (hence (20)) is classical; the contribution here is to (i) encode biologically motivated constraints (e.g., heterogeneous AT / GC composition, sequence context locality) as block invariance, (ii) measure distance to a set of steady states that preserves the initial block masses, and (iii) show that the minimal KL to this set remains a Lyapunov function even though the minimizing $\pi$ may change along the trajectory. This covers non-ergodic replication where global mixing is forbidden but within-block relaxation proceeds.
Remark 1 (Equality and strict decay).
Equality in (17) at step $n$ holds iff there exists a minimizer $\pi^{\star}\in\Pi(p_{0})$ with
$D_{\mathrm{KL}}\!\bigl{(}T_{j}p_{n}^{(j)}\,\|\,\pi_{j}^{\star}\bigr{)}=D_{\mathrm{KL}}\!\bigl{(}p_{n}^{(j)}\,\|\,\pi_{j}^{\star}\bigr{)}\quad\text{for all }j,$
i.e., DPI is tight blockwise (this typically means $p_{n}^{(j)}$ is already $T_{j}$-invariant). Otherwise the decrease is strict.
Corollary 1 (Per-step potential drop).
The drop $\Delta V_{n}:=V(p_{n})-V(p_{n+1})\geq 0$ for all $n$, with $\Delta V_{n}=0$ iff the condition in Remark 1 holds.
III Instantiations
III.1 Image replication via Gaussian convolution

### v2

# Kullback–Leibler Divergence Potential for Non-Ergodic Replication Dynamics: An Information-Theoretic Second Law

@@ Abstract
While previous research in information thermodynamics has focused on the thermodynamic costs associated with information “erasure” or “measurement” through concepts such as Landauer’s principle and mutual information, little systematic discussion has addressed the inherently irreversible nature of “replication” itself and the accompanying degradation of information structure. In this study, we construct a mathematical model of information replication using a discrete Markov model and Gaussian convolution, and quantify changes in information at each replication step: Shannon entropy, cross-entropy, and the Kullback–Leibler divergence (KLD). The monotonic decrease of the KLD potential $V$ (the minimal KLD to a reachable steady set) exhibits a Lyapunov-like property, which can be interpreted as a potential analogous to the free energy in the process by which a nonequilibrium system converges to a particular steady state. Furthermore, we extend this framework to the potential applicability to biological information processes such as DNA replication, showing that the free energy required for degradation and repair can be expressed in terms of KLD. This contributes to building a unified information-thermodynamic framework for operations such as replication, transmission, and repair of information.
1 General framework and main theorem
1.1 Block invariance and reachable steady set
[…]
These well-documented asymmetries motivate modeling replication on a block partition that separates AT- from GC-rich components. In our framework, block invariance captures slow exchange between AT and GC pools while allowing within-block relaxation; consequently, the KLD potential $V$ measures the departure from the stable block set under realistic composition-dependent mutation / repair biases. When such conserved coarse variables exist, the Markov dynamics does not converge to a single global invariant distribution; instead, the state space decomposes into components that evolve independently. We call this situation non-ergodic in the sense that long-term behavior retains memory of the initial coarse structure through the conserved weights.
@@ 1.1 Block invariance and reachable steady set
(i) In the image model, a blockwise Gaussian convolution forbids smoothing across block boundaries; the total intensity in each block is preserved at every step. (ii) In the DNA model, a block diagonal substitution kernel reflects biochemical or sequence constraints (e.g., AT- vs. GC-rich segments); block masses (AT vs. GC) remain fixed while within-block compositions relax. (iii) More generally, locality, topological separation, or routing restrictions in copying channels induce invariant ’blocks’ that prevent global mixing. Because replication preserves these coarse variables, the appropriate equilibrium notion is not a single point but a set of steady states obtained by fully relaxing within each invariant component while keeping the initial coarse weights fixed. This motivates modeling the state space as a disjoint union of blocks and defining a reachable steady set against which we will measure the distance using a KL-based potential. We emphasise that non-trivial block partitions $\mathcal{X}=\bigsqcup_{j=1}^{m}\mathcal{X}_{j}$ with $m>1$ are meaningful precisely when the dynamics itself possesses non-mixing components, so that probability mass cannot freely move between all microscopic states. In the fully mixing (ergodic) case, the only natural choice is the trivial partition with a single block, and our potential $V$ reduces to the standard KLD to a unique invariant steady state.
Framework and notation (matrix semantics made explicit)
We consider a finite state space partitioned into $m\in\mathbb{N}$ disjoint blocks
@@
with $|\mathcal{X}|=d<\infty$. The probability simplex over $\mathcal{X}$ is
$\Delta(\mathcal{X})\ :=\ \bigl\{\,p\in\mathbb{R}^{d}\ :\ p\geq 0,\ \mathbf{1}^{\top}p=1\,\bigr\},$
and we write $p=(p_{x})_{x\in\mathcal{X}}\in\Delta(\mathcal{X})$.
A one-step Markov kernel $T$ acts on the right by
$(pT)(y)=\sum_{x\in\mathcal{X}}p(x)\,T(x,y),\qquad p_{n+1}=p_{n}\,T,$
so $T$ is a $d\times d$ row-stochastic matrix ($T(x,y)\geq 0$ and $\sum_{y}T(x,y)=1$ for each row $x$). The matrix dimension $d\times d$ equals the number of states $|\mathcal{X}|$. Rows index the current state, columns the next state. Each block kernel $T_{j}$ has size $|\mathcal{X}_{j}|\times|\mathcal{X}_{j}|$ and updates the conditional $p^{(j)}$ while preserving its mass $w_{j}(p_{0})$. This row-vector/right-action convention is used throughout: $p_{n+1}=p_{n}\,T$.
Semantics: the row index is the current state, the column index is the next state, and the matrix size $d\times d$ equals the number of states. Restricting to a block $\mathcal{X}_{j}$ yields a subkernel $T_{j}$ of size $|\mathcal{X}_{j}|\times|\mathcal{X}_{j}|$.
Block-invariant dynamics.
The dynamics is block-invariant if the transitions never cross the block boundaries:
$x\in\mathcal{X}_{j}\ \Longrightarrow\ T(x,\mathcal{X}_{j})=1\qquad(j=1,\dots,m),$ (2)
equivalently $T(x,y)=0$ for $y\notin\mathcal{X}_{j}$ whenever $x\in\mathcal{X}_{j}$. Matrix-wise, $T$ is block diagonal: $T=\mathrm{diag}(T_{1},\dots,T_{m})$. Under (2), block masses are conserved and conditionals evolve block-wise:
$w_{j}(p_{n+1})=w_{j}(p_{n}),\qquad(Tp)^{(j)}=T_{j}\,p^{(j)}.$ (3)
Coarse variables (block masses) and within-block conditionals.
For $p\in\mathcal{P}(\mathcal{X})$ define
$w_{j}(p):=p(\mathcal{X}_{j})=\sum_{x\in\mathcal{X}_{j}}p(x),\qquad p^{(j)}(x):=\begin{cases}p(x)/w_{j}(p),&x\in\mathcal{X}_{j},\\ 0,&x\notin\mathcal{X}_{j}.\end{cases}$ (4)
Then $p$ decomposes as a convex mixture of conditionals within the block:
[…]
For each block, invariant measures and nonemptiness of the steady set is given by:
@@
Existence. Since every $T_{j}$ is a finite row-stochastic matrix, it admits at least one invariant distribution (e.g. by Brouwer’s fixed point theorem on the simplex or Perron–Frobenius for $T_{j}^{\top}$), hence $\mathcal{I}_{j}\neq\emptyset$ for all $j$. Given an initial distribution $p_{0}$, the reachable steady set is
$\Pi(p_{0}):=\Bigl\{\pi=\sum_{j=1}^{m}w_{j}(p_{0})\,\pi_{j}:\pi_{j}\in\mathcal{I}_{j}\Bigr\},$
which is nonempty. Intuitively, $\Pi(p_{0})$ collects all steady states that (i) fully relax within each block and (ii) preserve the initial coarse masses $w_{j}(p_{0})$ fixed by non-ergodic constraints.
Block primitivity.
Primitivity of a block kernel $T_{j}$ (irreducible and aperiodic; equivalently, $(T_{j})^{n}>0$ for some $n$) is invoked only to conclude that the within-block conditionals converge to a unique invariant $\pi_{j}^{\ast}\in\mathcal{I}_{j}$. For a block kernel $K$, with constant $\varepsilon$, means
$\exists\,\nu\in\Delta(\mathcal{X}_{j}),\ \varepsilon>0\ \text{ s.t. }\ K(x,\cdot)\ \geq\ \varepsilon\,\nu(\cdot)\quad\forall x\in\mathcal{X}_{j},$ (7)
which implies total-variation contraction at most $1-\varepsilon$ per step. We also define coefficient
$\delta(K)\ :=\ 1-\min_{x,x^{\prime}}\sum_{y\in\mathcal{X}_{j}}\min\{K(x,y),K(x^{\prime},y)\},$ (8)
which bounds the one-step total-variation contraction by $\delta(K)$.
1.2 Key Lemmas
@@
$(Tp)^{(j)}(A)$ $=\frac{(Tp)(A)}{w_{j}(p)}$
$=\frac{\sum_{x\in\mathcal{X}_{j}}p(x)\,T_{j}(x,A)}{w_{j}(p)}$
$=\sum_{x\in\mathcal{X}_{j}}\frac{p(x)}{w_{j}(p)}\,T_{j}(x,A)$
$=(T_{j}\,p^{(j)})(A).$ (11)
@@
Let $\pi=\sum_{j=1}^{m}w_{j}\,\pi_{j}$ with $w_{j}>0$ and $\pi_{j}\in\mathcal{P}(\mathcal{X}_{j})$ supported in disjoint blocks $\mathcal{X}_{j}$. Then, for any $p\in\mathcal{P}(\mathcal{X})$,
@@
Equations (9)–(10) show that the dynamics preserves coarse variables $\{w_{j}\}$ and evolves conditionals within each block, while (12) cleanly separates within-block divergences from a coarse-mass mismatch term. These identities will be the backbone for the Lyapunov property established in Sec. 1.3.
1.3 Lyapunov property of the KLD potential
Statement. Let $p_{n+1}=Tp_{n}$ with a block-invariant kernel $T=\mathrm{diag}(T_{1},\dots,T_{m})$ (Sec. 1.1).
@@ 1.3 Lyapunov property of the KLD potential
For all $n\geq 0$,
@@
Fix $\pi\in\Pi(p_{0})$ and write $\pi=\sum_{j=1}^{m}w_{j}(p_{0})\,\pi_{j}$ with $\pi_{j}\in\mathcal{I}_{j}$. Block masses are conserved along the trajectory (Lemma 1), hence $w_{j}(Tp)=w_{j}(p)=w_{j}(p_{0})$. By the KL block decomposition (Lemma 2),
@@
Applying the data-processing inequality (DPI) blockwise,
@@
and inserting (17) into (16) yields
@@
Taking the infimum over $\pi\in\Pi(p_{0})$ proves (14). If each $T_{j}$ is primitive with invariant $\pi_{j}^{\ast}$, then $p_{n}^{(j)}\to\pi_{j}^{\ast}$ and (15) follows. ∎
Robustness to weak inter-block leakage and re-partitioning
Motivation.
Secs. 1.21.3 establish a Lyapunov principle for $V$ under exact block invariance. To examine robustness to weak violations (small inter-block leakage or re-partitioning), we introduce a leakage-tolerant potential $V_{\delta}$ that continuously extends $V$ and preserves its qualitative behavior up to $O(\delta)$ slack.
Leakage-tolerant admissible set and closed form (corrected).
When leakage is small over the observation window, relax coarse masses by $\pm\delta$ and define
$\Pi_{\mathcal{P},\delta}(p_{0}):=\Bigl\{\,\pi\in\Delta(\mathcal{X})\ \Big|\ \forall j,\ \Bigl|\sum_{i\in\mathcal{X}_{j}}\pi_{i}-w_{j}(p_{0})\Bigr|\leq\delta\,\Bigr\},\qquad V_{\delta}(p):=\inf_{\pi\in\Pi_{\mathcal{P},\delta}(p_{0})}D_{\mathrm{KL}}(p\|\pi).$
By KL block decomposition and KKT optimality, the optimizer has the form
$w_{j}^{\star}\;=\;\biggl[\,\frac{w_{j}(p)}{\tau^{\star}}\,\biggr]_{[\,a_{j},\,b_{j}\,]},\quad a_{j}:=w_{j}(p_{0})-\delta,\ \ b_{j}:=w_{j}(p_{0})+\delta,$
where $\tau^{\star}>0$ is chosen so that $\sum_{j}w_{j}^{\star}=1$. Hence
$\boxed{\;V_{\delta}(p)=\sum_{j=1}^{m}w_{j}(p)\,\log\!\frac{w_{j}(p)}{\,w_{j}^{\star}\,},\quad w_{j}^{\star}=\biggl[\,\frac{w_{j}(p)}{\tau^{\star}}\,\biggr]_{[\,a_{j},\,b_{j}\,]},\ \ \sum_{j}w_{j}^{\star}=1.\;}$ (19)
Since $f(\tau):=\sum_{j}\bigl[w_{j}(p)/\tau\bigr]_{[a_{j},b_{j}]}$ is continuous and strictly decreasing in $\tau>0$, $\tau^{\star}$ is found by a short bisection on $(0,\infty)$.
A self-contained derivation and a sufficient threshold for strict decrease under leakage are provided in Appendix A.
Continuity and small-leakage persistence.
As $\delta\to 0$, $V_{\delta}(p)\to V(p)$. Hence, under sufficiently small leakage, the qualitative properties proved for $V$ (monotonicity along the dynamics, long-time limit, and the thermodynamic reading $k_{\mathrm{B}}T\,V_{\delta}$ per refresh) persist up to an $O(\delta)$ relaxation.
Coarsening by merging leaking blocks.
If mutually leaking blocks are merged to form a coarser partition $\mathcal{Q}\succeq\mathcal{P}$, data processing implies $V_{\mathcal{Q}}(p)\leq V_{\mathcal{P}}(p)$. Coarsening restores exact block invariance at the partition level, at the price of a more conservative (smaller) KL-based potential and a weaker maintenance-cost bound.
2 Instantiations
2.1 Image replication via Gaussian convolution
Block-patterned model.
First, we model an image with an explicit block (mosaic) structure: the pixel grid is partitioned into $m$ disjoint rectangular regions that remain isolated in the non-ergodic variant (no cross-region mass transfer). Figure 2 uses three blockings of the same $256\times 256$ image: $2\times 2$ (top), $4\times 4$ (middle), and $128\times 128$ (bottom). The ergodic variant corresponds to the same update without any blocking (global mixing).